OpenStudy (anonymous):

Square root of 45a to the fifth power

OpenStudy (dangerousjesse):

So \(\sqrt{45a}^5\)

OpenStudy (dangerousjesse):

Or \(\sqrt{45a^5}\)?

OpenStudy (anonymous):

The second

OpenStudy (dangerousjesse):

Okay

OpenStudy (dangerousjesse):

Hm.. Are you looking for the root, a plot, or what?

OpenStudy (anonymous):

The root

OpenStudy (imstuck):

The square root of 45 is equal to this:\[3\sqrt{5}\]and you can break up the a^5 like this:\[\sqrt{a ^{4}\times a ^{1}}\]and that a^4 comes out as an a^2 so in the end you have this:\[3a ^{2}\sqrt{5a}\]

OpenStudy (anonymous):

Thank you so much no i can go to bed

OpenStudy (imstuck):

Oh yay! Go to bed! Good for you!

OpenStudy (imstuck):

TY for the medal, btw!